OpenStudy (anonymous):

Form a polynomial whose zeros and degree are given. Zeros: 2, multiplicity 2; 3, multiplicity 1; degree 3 A. x^3+7x^2+16x+12 B. x^3+7x^2+12x+12 C. x^3-4x^2+16x-12 D. x^3-7x^2+16x-12

OpenStudy (anonymous):

So we're given that are zeros/roots are: 2 with a multiplicity 2 3 with a multiplicity 1 In other words, our roots are 2, 2, 3! Working backwards, we see that if those are the roots, then that means that x=2, x=2, and x=3 x-2=0, x-2=0, x-3=0 (x-2)(x-2)(x-3)=0 Now multiply those terms and you'll get your answer!

OpenStudy (anonymous):

What terms do you multiply??

OpenStudy (anonymous):

(x-2)(x-2)(x-3)

OpenStudy (anonymous):

so the answer would be, x^3-3x^2+4x-12?

OpenStudy (anonymous):

But that's not one of my answers...

OpenStudy (anonymous):

You're probably made a mistake when you were solving algebraically. (x-2)(x-2)(x-3) (x-2)^2 (x-3) (x^2-4x+4)(x-3) Is this what I have so far? Do you understand how I got there?

OpenStudy (anonymous):

you*

OpenStudy (anonymous):

yes i got that.